The potential at a distance r from a positive point charge is +V. What is the potential at a distance of r/3? 0 +3V A charge o +20 and -2Q are located at two of the vertices of an equilateral triangle. Which of the following is true? O The potential at the triangles empty vertex is positive. O The potential at the triangles empty vertex is negative. O The potential at the triangles empty vertex is zero.
Answer #1

a)
Consider q as the positive charge.
V = kq/r
Where k = 1/4\pi\epsilono

When r becomes r/3,
New potential, V' = kq/(r/3)
= 3kq/r
= + 3V

b)
Consider the length of a side as 'a'
Potential at the empty vertex = k(+2Q)/a + k(-2Q)/a
= 2kQ/a - 2kQ/a
= 0
i.e, the potential at the empty vertex is zero.